# Understanding Self-Regulated Learning Behavior Among High and Low Dropout Risk Students During CS1: Combining Trace Logs, Dropout Prediction and Self-Reports

### v1

@@ 4. Results
4.1. Trace log preprocessing and learning tactics
@@
Figure 3. Cluster validation for clustering of student sessions. Cluster validation indices are rescaled to fit the figure and adjusted so that higher values are better.
Code N Description Example codes in sessions
F_CourseMat_Examples 509 Focusing on course materials and examples in the materials
answer:materials-book-example
read:materials-book
F_Lec_Engaged 766 Focusing on the lecture and lecture materials
read:lecture-cur-wk
answer:lecture-cur-wk-example
F_Lec_Video 469 Focusing on lecture videos
read:lecture-cur-wk
watch-video:lecture-cur-wk
F_CurTasks_Intro 350 Focusing on current week’s introductory tasks
answer-wrong:tasks-cur-intro
answer:tasks-cur-intro
F_CurTasks_Core_Attempt 356 Focusing on current week’s core tasks, submitting incomplete solutions
answer-wrong:tasks-cur-core
read:tasks-cur
F_CurTasks_Core_Correct 410 Focusing on current week’s core tasks, submitting correct solutions
answer:tasks-cur-core
read:tasks-cur
F_CurTasks_Basic_Attempt 369 Focusing on current week’s basic tasks, submitting incomplete solutions
answer-wrong:tasks-cur-basic
read:tasks-cur
F_CurTasks_Basic_Correct 529 Focusing on current week’s basic tasks, submitting correct solutions
read:tasks-cur
answer:tasks-cur-basic
F_CurTasks_Extra 640 Focusing on current week’s extra tasks (bonus, guru) or supplementary tasks
read:materials-book
read:tasks-extra
TA_Sess 57 Attending one-to-one session with TAs
session-start
session-end
F_PrevTasks_Basic 361 Focusing on reviewing last week’s tasks, focusing on fixing basic answers
read:tasks-prev
answer:tasks-prev-basic
answer-wrong:tasks-prev-basic
F_PrevTasks_Deep 629 Focusing on reviewing last week’s tasks, reviewing and fixing multiple tasks
read:tasks-prev
answer:tasks-prev-core
answer:tasks-prev-intro
Table 1. Weekly learning tactics interpreted from student session clusters. For each tactic, short code, number of exemplary student sessions and description of the tactic is given.
Overall, the trace logs were split into 7,226 individual student sessions. After removing sessions with just one event, 5,445 sessions were included in the analysis. To select a reasonable number of session clusters, multiple cluster validation indices (CVI) were computed. The clustering performance of the $k$-medoids clustering on the student sessions is depicted in Figure 3. Given the weekly schedule of the course (Figure 1), we expected to select at least 7 session types. We ultimately selected $K_{\text{sess}}=12$ session clusters, as the CVIs for this cluster count are close to best and because a larger number of clusters would allow capturing more detail in students’ weekly strategies. From the session clusters, we interpreted 12 tactics the participants used during CS1. The descriptions of all tactics, the number of student sessions and some example event codes in the student sessions are listed in Table 1.
@@
It must be noted, that while the table shows distinct event codes for each tactic, in practice student sessions may contain a mix of similar events. For example, both “Focusing on current week’s basic tasks, submitting correct solutions” and “Focusing on current week’s basic tasks, submitting incomplete solutions” tactics have examples of students answering wrongly (answer-wrong:tasks-cur-basic) and correctly (answer:tasks-cur-basic), but the proportion of the correct answers in the sessions is much higher for the former tactic (32% answering correct vs. 7% answering wrong) than the latter tactic (10% answering correct vs. 37% answering wrong). Therefore, the relative proportions of recorded event types within each cluster were primarily used to interpret the student session clusters as tactics.
4.2. Weekly learning strategy types
@@
Figure 4. Cluster validation for clustering of the weekly learning strategies. Cluster validation indices are rescaled to fit the figure and adjusted so that higher values are better.
@@
Figure 5. The identified weekly strategy clusters; each cluster shows multiple weekly learning strategies as vertical bars; colors in weekly learning strategies indicate learning tactics listed in the figure legend and described in more detail in Table 1; the size of each area describes the relative frequency of the used tactic within the strategy; for each cluster, interpreted strategy type, number of strategies in the cluster $N_{\text{strat}}$ and average predicted dropout probability $\bar{p}_{\text{drop}}$ is given.
(a) Heuristic Net of strategies in the cluster that was identified as Falling behind, realizing struggle too late.
(b) SHAP waterfall plot explaining prediction for a weekly learning strategy of a single student inside the cluster that was identified as Falling behind, realizing struggle too late.
Figure 6. Example of graphs used to identify learning strategy type Falling behind, realizing struggle too late.
After recoding student sessions using the learning tactics and splitting sessions by course week, we ended up with 458 weekly learning strategies for further analysis. After applying weekly dropout predictions to group weekly learning strategies, 396 strategies were categorized as low dropout risk strategies ($p_{\text{drop}}\leq 0.5$) and 62 as high dropout risk strategies ($p_{\text{drop}}>0.5$). Initial cluster validation for the Expectation Maximization algorithm using BIC suggested using a diagonal covariance matrix. Based on further evaluation with multiple CVIs (Figure 4), we selected $K_{\text{pass}}=3$ clusters for low dropout risk strategies and $K_{\text{drop}}=9$ clusters for high dropout risk strategies. Here, we preferred to select a larger cluster number to better capture various learning strategy types. To identify the learning strategy types, we plotted the frequencies of the learning tactics used in every weekly strategy, resulting in Figure 5. In addition, we generated heuristic nets for each cluster to aid interpretation. In some edge cases, we also reviewed the predictions of the used dropout prediction model using SHAP waterfall plots. Examples of both are provided in Figure 6. Overall, we interpreted the following learning strategy types from clusters of low dropout risk strategies:
@@ 4.3. Student learning profiles
(b) Dendrogram of the clustered student profiles. The vertical axis includes identification numbers of each student. The vertical line indicates the selected cluster count $K_{\text{student}}=5$.
@@
Figure 7. Cluster validation for student profile clusters using CVIs and a visual approach using a dendrogram.
@@
Figure 8. Student profile clusters displayed as the number of students that used each weekly strategy type throughout all course weeks. For each cluster, the interpreted profile type name, number of students in the cluster $N_{\text{student}}$ median percentage of the completed weekly tasks $\text{Md}(w)$, median percentage of exam points $\text{Md}(w)$, median course grade on scale 0–5 $\text{Md}(g)$ and mean dropout prediction over all weekly strategies $\bar{p}_{\text{drop}}$, is listed. For clusters with students who filled out the self-report, themes mentioned by at least half of the answerers are listed (see Appendix C for a list of all themes and descriptions). Additionally, a distribution of self-reported opinions of own effort are shown. Cluster interpreted as Course dropouts does not include any students with self-reports.
@@
We used CVIs and a dendrogram to select the linkage criterion for clustering student profiles. Initial validation suggested to use complete linkage criterion. In further validation for cluster numbers, multiple CVIs (Figure 7(a)) suggested 4–6 student profile clusters. Based on a visual review of a resulting dendrogram (Figure 7(b)), one student’s profile (Student 1) appears to strongly differ from the other students. As such, we ultimately selected $K_{\text{student}}=5$ as the cluster number, with four clusters representing learning profile types and one cluster containing the outlier student.
@@
An overview of the interpreted learning profile types along with a summary of the available self-report data is shown in Figure 8. A further listing and description of all emerged themes used in the study’s analysis is listed in Appendix C. Overall, three of the learning profiles represent various levels of activity during CS1, while one learning profile includes all students that dropped out mid-course. Further, the profile of the single outlier student represents a case of a low-activity student who managed to pass the course by compensating low task performance with good exam results. All clusters include students without self-reports, with the cluster of course dropouts having no recorded self-reports at all. The profile of a single student contains a self-report that was extensively used in interpreting the profile. Nevertheless, visually, there appear to be some differences in self-reported opinions of students’ own performance between the clusters. Next, we describe the learning profile types interpreted from the clusters based on the visualizations, descriptive statistics, confirmatory tests, and excerpts from the self-reports.
@@ 4.3.1. Adaptive understanding seekers
The first interpreted profile type ($N_{\text{student}}=20$) contains students who primarily sought understanding and adapted their study strategies as needed. Generally these students have stuck to the understanding-seeking strategy throughout the course. However, some students do appear to have varied their strategies during the course to adjust to new topics or find other ways of learning, as exemplified in the quotes below:
@@
Some students occasionally slipped into high dropout risk strategies, but eventually managed to recover. Based on the self-reports, many students in this profile type used various resources, including external sources from the internet while remaining critical. An example is provided via a quote below:
@@
As such, the profile represents students who really wanted to ensure understanding of the topic while progressing through the course. It must be noted that not all understanding seekers succeeded with high grades; those who self-reported bad experiences with the course attributed it to external factors such as course workload and personal life:
[Student 19]: The course has a high workload and for me required much more effort than 10–12 hours weekly.
@@
[…]
@@ 4.3.2. Diligent course content followers
The second profile type ($N_{\text{student}}=14$) represents students who conscientiously used the course materials and appeared to closely mimic the suggested weekly schedule (cf. Figure 1). Their primary strategy was material-focused during the majority of the course duration. However, toward the end of the course, more students shifted to the understanding-seeking strategy. This possibly indicates some students’ need to enhance comprehension before final exams as the example course schedule may not have been enough to gain sufficient understanding of the topics. In self-reports, students emphasized the importance of seeking help:
@@
Students also reported relying heavily on the teaching assistants, implying some dependence to complete tasks and even to self-regulate:
@@
[…]

### v2

@@ 5. Results
We present the results of each step of the analysis in the order of the pipeline depicted in Figure 2 and described in Section 4.
5.1. Log preprocessing
The raw trace log consisted of 86,738 events in total. After splitting the events using the 25-minute cut-off rule, the log was divided into 7,226 individual student sessions. After removing sessions with just one event, 5,445 student sessions were included in the subsequent analysis.
5.2. Learning tactic discovery
@@
Figure 4. Cluster validation for clustering of student sessions. Cluster validation indices are rescaled to fit the figure and adjusted so that higher values are better. The selected number of clusters ($K_{\text{sess}}=12$) is indicated with a dashed line.
To select a reasonable number of clusters for grouping raw events into learning tactics, we computed multiple cluster validation indices (CVIs). The clustering performance of the $k$-medoids clustering on the student sessions is depicted in Figure 4. Based on the figure, three CVIs suggest 7 clusters (Medoid Silhouette, Xie-Beni and Dunn’s Index). However, a slight elbow appears at 12 clusters for several CVIs (S Dbw, Davies-Bouldin, Xie-Beni), suggesting a potentially interesting alternative clustering.
Based on the weekly schedule of the course (see Figure 1 and Section 3.1), we expected to select at least 7 session types related to reading course materials, attending lectures, completing different types of tasks (introductory, core, basic, bonus/guru), and reviewing previous weeks’ tasks. We were also open to selecting a more fine-grained clustering to uncover more specific patterns in students’ learning activities. Being familiar with the course context, we therefore reviewed the session clusters for both 7 and 12 clusters. We ultimately selected $K_{\text{sess}}=12$ session clusters, as this clustering captured more detailed activities. Importantly, unlike the 7-cluster solution, the 12-cluster solution captured one-on-one TA sessions as a separate tactic.
From the session clusters, we interpreted 12 tactics that the participants used during CS1. We referred to the frequency of each event code in each cluster to form our interpretation. The event codes and frequencies used for interpretation are provided in detail in Appendix C. Overall, we identified the following learning tactics:
  • Focusing on course materials and examples (F_CourseMat_Examples, $N=509$): This tactic is characterized by a focus on the course materials, such as the course book. Students read the course book and interacted with the programming examples provided in it. These are mostly runnable code examples that students were free to adjust, but they also include interactive simulators for visualizing abstract concepts, such as memory references.
  • Focusing on the lecture and lecture materials (F_Lec_Engaged, $N=766$): In this tactic, a student engaged actively with the lecture materials for a specific course week. The tactic includes events related to attending the lecture, although the relative frequency of these events is lower than that of interacting with the lecture materials. The lecture materials included additional interactive examples and questions intended for use during the lecture; compared to simply watching lecture videos, this tactic is characterized by higher interaction with the lecture examples and with questions intended for the live lecture.
  • Focusing on lecture videos (F_Lec_Video, $N=469$): Compared with the previous tactic, this tactic shows a higher relative frequency of watching lecture videos and a lower relative frequency of interacting with the lecture materials. We therefore characterize it as a student watching lecture videos, either after the lecture to revisit it or in place of the original live lecture. This tactic also captures events related to viewing the previous week’s lecture materials, further emphasizing a focus on the lecture contents after the live lecture.
  • Focusing on the current week’s introductory tasks (F_CurTasks_Intro, $N=350$): This tactic relates to focusing on introductory tasks for the current week. Since the introductory tasks are not mandatory for passing and are meant as a scaffolding tool, the number of sessions in this cluster is lower than for tactics focusing on basic or core tasks.
  • Focusing on the current week’s core or basic tasks, attempts and correct submissions: The focus on the course’s core and basic tasks is split into four separate tactics. The first two relate to core tasks: the attempting tactic (F_CurTasks_Core_Attempt, $N=356$) is characterized by a high number of incorrect submissions of core tasks, suggesting a tactic in which a student either tests the tasks, is in the process of solving them, or solves them only after several initial attempts; the correct submissions tactic (F_CurTasks_Core_Correct, $N=410$) in turn contains a high number of correct submissions of core tasks, suggesting a tactic in which a student either quickly solves the core tasks after an initial incorrect attempt or solves them after multiple attempts. Similarly, two tactics exist for basic tasks: one for attempting (F_CurTasks_Basic_Attempt, $N=369$) and one for correct submissions (F_CurTasks_Basic_Correct, $N=529$).
  • Focusing on the current week’s extra or supplementary tasks (F_CurTasks_Extra, $N=640$): This tactic captures students’ focus on materials beyond the course book and mandatory exercises. These materials are primarily aimed at supporting students with the core tasks, for example through simplified guided exercises. The tactic also includes sessions focused on advanced bonus or guru tasks, but the relative frequency of these events is small.
  • Attending a one-on-one session with a TA (TA_Sess, $N=57$): This tactic is characterized by attending one-on-one sessions with TAs as described in Section 3.1. While the primary events of this tactic are starting and ending the session, the tactic also includes a small number of other events, suggesting that the TA and the student occasionally looked at course materials beyond the course mini-project.
  • Reviewing and fixing previous week’s basic answers (F_PrevTasks_Basic, $N=361$): In this tactic, a student reviews and fixes answers from the previous week. It primarily includes submitting answers to the previous week’s basic questions, watching the task review session videos and viewing model answers. Overall, the tactic suggests a basic review of the previous week’s tasks, inspecting the student’s own answers and trying out model answers.
  • Deeply reviewing and fixing the previous week’s tasks (F_PrevTasks_Deep, $N=629$): The final identified tactic relates to reviewing and likely fixing answers to the previous week’s tasks. Given that students were required to review and fix their wrong answers in this course (see Section 3.1), this tactic includes a variety of events related to submitting answers to various old tasks.
@@
We note that while each tactic is associated with distinct event codes, in practice student sessions may contain a mix of similar events. For example, both the correct submissions and attempting tactics for basic tasks include events where students answered wrongly (answer-wrong:tasks-cur-basic) and correctly (answer:tasks-cur-basic), but the proportion of correct answers in the sessions is much higher for the former tactic (32% correct vs. 7% wrong) than for the latter (10% correct vs. 37% wrong). As such, the interpretation focused only on the most prominent event codes in each cluster, with less frequent events treated as noise.
5.3. Learning strategy formulation
Using the above clusters, we recoded each student session into its associated learning tactic and extracted the timestamp of the first event in each session. We then grouped the sessions by student, and for each student’s sessions we further grouped them by course week using the time of the first event. This yielded 458 weekly learning strategies in total.
5.4. Learning strategy characterization
@@
Figure 5. Cluster validation for clustering of the weekly learning strategies. Cluster validation indices are rescaled to fit the figure and adjusted so that higher values are better. The selected number of clusters ($K_{\text{drop}}=3$ and $K_{\text{pass}}=9$) are indicated with dashed lines.
@@
Figure 6. Summary of the identified weekly strategy clusters. For each cluster, the interpreted strategy type, the number of strategies in the cluster $N_{\text{strat}}$, and the average predicted dropout probability $\bar{p}_{\text{drop}}$ are given. Each cluster includes a heuristic net that depicts the most important transitions between tactics within the cluster. The relative frequencies of all learning tactics in the cluster are shown as bar plots; the colors representing the learning tactics are listed in the figure legend and described in Section 5.2.
Using the weekly dropout predictor described in Section 4.4 for each student, we grouped the weekly learning strategy logs into low- and high-dropout risk strategies. In total, 396 strategies were categorized as low dropout risk ($p_{\text{drop}}\leq 0.5$) and 62 as high dropout risk ($p_{\text{drop}}>0.5$).
As with the identification of learning tactics, we used cluster analysis with multiple CVIs to explore the types of weekly study behavior. Initial cluster validation for the EM clustering algorithm using BIC suggested using a diagonal covariance matrix. The clustering performance of the EM algorithm on low- and high-dropout risk strategies is presented in Figure 5. For low-risk strategies, three CVIs suggest 10 clusters. However, these indices appear to exhibit a linear monotonic increase after $K=8$ while the other three CVI values decrease, suggesting possible over-segmentation rather than true structural discovery (Liu et al., 2010). Because we had no prior expectations about the number of behavior types we expected to see, we aimed to select the highest cluster number supported by multiple CVIs in order to obtain as fine-grained a clustering as possible while avoiding over-segmentation. Overall, we selected $K_{\text{pass}}=3$ clusters for low dropout risk strategies, as two CVIs attain the highest value there (Dunn’s Index, S Dbw) and it is a notable elbow point for two other CVIs (Xie-Beni, Davies-Bouldin). Similarly, we selected $K_{\text{drop}}=9$ clusters for high dropout risk strategies, as multiple CVIs attain a high value and it is an elbow point for some of the CVIs.
Figure 6 summarizes the clusters of low- and high-dropout risk strategies. For each cluster, we include the distribution of learning tactics. We also generated heuristic nets for each cluster to aid interpretation. The nets were heuristically simplified to include only transitions between the most frequent events. When interpreting the clusters, we also referred to dropout predictions and performance data (see Section 4.4 for a description of the available data).
Overall, we interpreted the following learning strategy types from clusters of low dropout risk strategies:
@@ 5.5. Student characterization
(b) Dendrogram of the clustered student profiles. The vertical axis includes identification numbers of each student.
@@
Figure 7. Cluster validation for student profile clusters using CVIs and a visual approach using a dendrogram. The selected number of clusters ($K_{\text{student}}=5$) is indicated with a dashed line.
@@
Figure 8. Student profile clusters displayed as the number of students that used each weekly strategy type throughout all course weeks. For each cluster, the interpreted profile type name, number of students in the cluster $N_{\text{student}}$ median percentage of the completed weekly tasks $\text{Md}(w)$, median percentage of exam points $\text{Md}(w)$, median course grade on scale 0–5 $\text{Md}(g)$ and mean dropout prediction over all weekly strategies $\bar{p}_{\text{drop}}$, is listed. For clusters with students who filled out the self-report, themes mentioned by at least half of the answerers are listed (see Appendix D for a list of all themes and descriptions). Additionally, a distribution of self-reported opinions of own effort are shown. Cluster interpreted as Course dropouts does not include any students with self-reports.
@@
Initial validation suggested using the complete linkage criterion in agglomerative clustering. To select the appropriate number of student profile clusters, we used both CVIs (Figure 7(a)) and a dendrogram (Figure 7(b)). On closer inspection of the CVIs, $K=4$ and $K=5$ emerged as potential cluster numbers, as some CVIs either reach the maximum value or show a clear increase. Based on an additional visual review of the dendrogram, one student’s profile appears clearly separate from the others. We therefore selected $K_{\text{student}}=5$ as the cluster number, with four clusters representing learning profile types and one cluster containing the outlier student.
@@
An overview of the interpreted learning profile types, along with a summary of the available self-report data, is shown in Figure 8. A full listing and description of all themes that emerged in the study’s analysis is provided in Appendix D. Three of the learning profiles represent various levels of activity during CS1, while one learning profile includes all students who dropped out mid-course. The profile of the single outlier student represents a low-activity student who managed to pass the course by compensating for low task performance with good exam results. All clusters include some participants without submitted self-reports, and the cluster of course dropouts contains no self-reports at all. The profile of the single outlier contains one self-report that was used extensively to characterize the student. Nevertheless, visually, there appear to be some differences between clusters in the self-reported opinions of students’ own performance. Next, we describe the learning profile types interpreted from the clusters based on the visualizations, descriptive statistics, confirmatory tests, and excerpts from the self-reports. When reporting the excerpts, we use student identifiers in relation to all original participants in the study ($N=47$). The same identifiers appear in Figure 7(b).
@@ 5.5.1. Adaptive understanding seekers
The first interpreted profile type ($N_{\text{student}}=20$) contains students who primarily sought understanding and adapted their study strategies as needed. In general, these students stuck to the understanding-seeking strategy throughout the course, but they also adjusted their strategies on a weekly basis. Based on the available reports (13 students), most of the adjustment was made in response to personal circumstances, such as switching from on-site lectures to online videos owing to a lack of time or changes in personal life. Some students, however, adjusted their strategies based on reflections on their learning, as exemplified in the quotes below:
@@
[Student 4]: I noticed that theory-focused learning is more natural and familiar to me […] I made a lot of notes to the tasks and course book. […] Compared to my previous experiences I changed my study methods in a way by using more time on exercises, as this is clearly the best way to learn…
Nevertheless, most of the self-reports in this cluster (11 out of 13) emphasized the importance of developing regular study habits, indicating that the adaptation was mostly a response to the weekly topic or changed circumstances. Based on the learning strategies, some students occasionally slipped into high dropout risk strategies but eventually managed to recover by adjusting their strategy. However, there is no mention of this in the self-reports.
All of the students with self-reports in this cluster reported using multiple learning resources, with the course book (12 reported), lectures (10 reported) and weekly group learning sessions (9 reported) being the most commonly used. Some students also used materials external to the course, as illustrated in the quote below:
@@
As such, this cluster represents students who really wanted to ensure understanding of the topic while progressing through the course. It must be noted that not all students in this cluster succeeded with high grades; those who self-reported bad experiences with the course attributed it to external factors such as course workload and personal life:
[Student 19]: The course has a high workload and for me required much more effort than 10–12 hours weekly. I enrolled in the course out of interest, which is why other studies took priority.
@@
While the self-reported opinion of own progress appears visually positive, the Brunner-Munzel test suggests that the opinions in this profile type’s cluster are stochastically equal to the opinions of other students who answered the self-report ($\hat{p}^{*}(25.886)=0.283,p=.779$). In other words, the opinions of students in this profile type and of the other students tended to be the same. Similarly, the final course grades of this profile type’s students and other study participants are stochastically equal ($\hat{p}^{*}(44.684)=-0.953,p=.346$). On the other hand, the dropout predictions for the weekly strategies used by students of this profile type and those of the other study participants are not stochastically equal ($\hat{p}^{*}(434.690)=-3.351,p<.001,\hat{p}^{\prime\prime}=0.589$); the effect size indicates that students of this profile type tended to use strategies with lower dropout probability than other participants.
@@ 5.5.2. Diligent course content followers
The second profile type ($N_{\text{student}}=14$) represents students who conscientiously used the course materials and appeared to follow the suggested weekly schedule described in Section 3.1. Based on the frequency of the learning strategies (Figure 8), their primary strategy was material-focused for most of the course. However, toward the end of the course, more students shifted to the understanding-seeking strategy. This may indicate a need to strengthen comprehension before the final exam, as the suggested course schedule may not have been enough to gain a sufficient understanding of the topics. All students with self-reports in this profile type (12 students) reported following the course schedule, with most attending lectures on-site and some watching lecture recordings. Only one student reported skipping lectures at the start of the course:
[Student 44]: At the beginning, I attended lectures, but right during the first weeks of the course I realized that I absolutely cannot concentrate there and that it was a waste of time. After that, I went directly to the lecture examples and to the weekly exercises.
All students with self-reports in this profile type also reported relying on some form of help-seeking resource, with weekly group learning sessions (7 reported), peer support groups (7 reported) and friends or relatives (6 reported) being the most commonly mentioned. Looking more closely at the reports, most students emphasized the importance of seeking help:
@@
Some students also reported relying heavily on the teaching assistants, suggesting some dependence on them to complete tasks and even to self-regulate:
@@
In this profile, the self-reported opinions are visually more negative than in other profiles. However, similar to adaptive understanding seekers, the opinions of students of this profile type appear to be stochastically equal to opinions of other answerers ($\hat{p}^{*}(17.963)=-0.481,p=.637$). Further, the final course grade is also stochastically equal to that of other students ($\hat{p}^{*}(39.376)=-1.297,p=.202$). In turn, predicted dropout probability of weekly strategies that the students of this profile type used is not stochastically equal to those of other study participants ($\hat{p}^{*}(453.337)=2.143,p=.033,\hat{p}^{\prime\prime}=0.443$). The effect size suggests the students in this profile tended to use strategies with higher dropout probability than other students.